OpenStudy (anonymous):

Show that the set X with the given operations fails to be a vector space by identifying all axioms that hold and fail to hold: The set X=R^3 with vector addition + defined by: (a,b,c)+(x,y,z)=(1,y,c+z) and scalar multiplication . defined by: k . (a,b,c)=(ka,kb,kc)

OpenStudy (zarkon):

no zero vector for addition

OpenStudy (zarkon):

there should be a vector, \(\vec{0}\), such that \[<a,b,c>+\vec{0}=<a,b,c>\]

OpenStudy (turingtest):

Correct me if I'm wrong Zarkon, but we also have no \(-\vec u\) such that \(\vec u+(-\vec u)=\vec0\) , right? also we don't have \((c+k)\vec u=c\vec u+k\vec u\) because of the definition of addition here so this is not a vector space for quite a few reasons

OpenStudy (zarkon):

Correct...I just gave one reason why it was not a vector space.

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

Consider the vector space M22 of all 2*2 matrices. 2.1 Show that B ={A1,A2,A3,A4} is a basis for M22 where: A1[3 6]; A2=[0 -1]; A3= [0 -8]; and A4= [1 0] [3 -6] [-1 0] [-12 -4] [-1 2] 2.2 Write A=[6 2] [5 3] 2.3 Prove that the subset D={A element of M22: A^T+A=0} forms a subspace of M22